solve the system by substitution

8x-4y=16
y=2x-4

Answer by tutorcecilia(2152)

y=2x-4 [substitute for the y-term in the first equation]
8x-4y=16
8x-4(2x-4)=16 [simplify]
8x-8x+16=16
0+16=16
16=16 [Infinate number of solutions because the equations represent the same line]